# Separable commutative algebras in equivariant homotopy theory

### v1

@@ Abstract.
Given a finite group $G$ and a commutative ring $G$-spectrum $R$, we study the separable commutative algebras in the category of compact $R$-modules. We isolate three conditions on the geometric fixed points of $R$ which ensure that every separable commutative algebra is standard, i.e. arises from a finite $G$-set. In particular we show that all separable commutative algebras in the categories of compact objects in $G$-spectra and in derived $G$-Mackey functors are standard provided that $G$ is a $p$-group. In these categories we also show that for a general finite group $G$, we cannot expect all separable commutative algebras to be standard.
@@ 1. Introduction
The study of separable commutative algebras in the setting of tensor-triangular geometry was initiated by Balmer in [Bal11]. Separable commutative algebras are useful in this context for several reasons. Firstly, they allow to form categories of modules internally to the world of triangulated categories, without the necessity of assuming that the triangulated category admits a model. Building on this, Balmer showed that restriction functors in (modular) representation theory can be understood as an extension-of-scalars functors [Bal15], and proved a generalization of the Neeman-Thomason Localization Theorem [Bal16a]. Secondly, separable commutative algebras allow us to formulate a going-up theorem in tt-geometry [Bal16b], which is a useful tool for calculating the Balmer spectrum of a triangulated category, and powerful descent results in tt-geometry [Bal16b],[BCH+24] and [BCH+23]. Finally, separable commutative algebras are tightly related to the étale topology of the given tt-category. A result of Neeman [Nee18] informally says that the only separable commutative algebras in the perfect derived category of a Noetherian scheme are those arising from finite étale morphisms. This connection with the étale topology was strengthened even further by the first two authors in [NP24] where they showed that separable commutative algebras recover finite étale algebras in various contexts where a notion of étaleness already exists.
@@
Informally this problem ask whether the étale topology which appears in modular representation theory is richer than what is produced by subgroups. A positive answer to this question was given by Balmer-Carlson [BC18] in the case of cyclic $p$-groups and by the first two authors [NP24] for $p$-groups of $p$-rank one. This question is completely open outside of these cases.
@@
    the commutative algebras $(\Phi^{K}(R))^{hW_{G}(K)}$ and $(\Phi^{K}(R))^{tW_{G}(K)}$ are indecomposable.
@@
    If $\Phi^{K}(R)$ is a retract in $\operatorname{Mod}(\Phi^{K}(R))^{tW_{G}(K)}$ of the coinduced algebra $(\Phi^{K}(R))^{W_{G}(K)/U}$ for some $U\subseteq W_{G}(K)$, then $U=W_{G}(K)$, see 6.1.
@@
We will show that the above conditions are satisfied if $R=\mathbb{S}_{G}$ is the sphere $G$-spectrum, and also if $R=\operatorname{infl}_{e}^{G}\mathbb{Z}$, provided that $G$ is a nontrivial $p$-group. The latter case has gained increasing interest for its connection to the category of derived $G$-Mackey functors of Kaledin, see [PSW22]. Our main result then states:
Theorem (9.6).
@@
is an equivalence.
It follows that if $G$ is a $p$-group, then all separable commutative algebras are standard, both in compact $G$-spectra and in compact derived $G$-Mackey functors. As an immediate application, we also show that the Galois groups of these categories, in the sense of [Mat16], are trivial, see 9.8. The techniques developed in this paper also allow us to produce some counterexamples. Indeed we show in Section 10 that if $G=C_{6}$, then there are separable commutative algebras in the above two categories which are not standard.
@@ Acknowledgements
We thank Bastiaan Cnossen, John Greenlees and Achim Krause for useful conversations.
The first and second authors are supported by the SFB 1085 Higher Invariants in Regensburg. The third author was supported by the Danish National Research Foundation through the Copenhagen Centre for Geometry and Topology (DNRF151) by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 427320536 - SFB 1442, as well as under Germany’s Excellence Strategy EXC 2044 390685587, Mathematics Münster: Dynamics-Geometry-Structure. .
@@ 3. Homotopy $G$-fixed points and separable algebras
Give a group $G$ and $\mathscr{C}\in\operatorname{Fun}(BG,\operatorname{Cat}_{\infty})$, we can form
@@
where $K^{g}=gKg^{-1}$, and the restriction from $K$ to $H\cap K^{g}$ is along the conjugation map $H\cap K^{g}\subset K^{g}\cong K$. This is a consequence of [CLL23, Lemma 4.9] with $S=T=BG$ and $\mathcal{C}$ the $BG$-category given by
[…]
@@
${\coprod_{g\in H\backslash G/K}G/(H\cap K^{g})}$${BH}$${BK}$${BG}$
is an equivalence. One also checks that the limit-preserving extension of ()(*)( ∗ ) to $\mathscr{S}_{/BG}$ sends $BH\to BG$ to $\operatorname{Mod}(A)^{hH}$.
@@
[…]
@@ 4. Stable module categories and the Tate construction
${X}$${X}$${X/ab}$${X}$${X}$${X/a}$$\scriptstyle{b}$$\scriptstyle{ab}$$\scriptstyle{\mathrm{id}}$$\scriptstyle{a}$
@@
00{0}${(\pi_{*}(X))/ab}$${\pi_{*}(X/ab)}$${(\pi_{*}(\Sigma X))[ab]}$00{0}00{0}${(\pi_{*}(X))/a}$${\pi_{*}(X/a)}$${(\pi_{*}(\Sigma X))[a]}$${0.}$$\scriptstyle{\mathrm{red}}$$\scriptstyle{\cdot b}$
@@
00{0}${(\pi_{*}(X))/ab}$${\pi_{*}(X/ab)}$${(\pi_{*}(\Sigma X))[ab]}$00{0}00{0}${(\pi_{*}(X))/a\oplus(\pi_{*}(X))/b}$${\pi_{*}(X/a)\oplus\pi_{*}(X/b)}$${(\pi_{*}(\Sigma X))[a]\oplus(\pi_{*}(\Sigma X))[b]}$${0.}$$\scriptstyle{\mathrm{red}}$$\scriptstyle{(\cdot b,\cdot a)}$
@@
Since cofibres in $\operatorname{Cat}_{\infty}^{\mathrm{perf}}$ are given by the idempotent completion of the Verdier quotient, we need to check that the image of $\mathrm{Nm}_{G}$ is $\operatorname{thick}_{\otimes}\langle A^{G/e}\rangle$. By [CMNN24, Proposition 2.18] taking $\mathcal{A}=\operatorname{Perf}(A)\in\operatorname{Fun}(BG,\operatorname{Cat}_{\infty}^{\mathrm{perf}})$ there, we know that the image of the norm map agrees with $\operatorname{thick}(A^{G/e})$ which is automatically an ideal by an application of the projection formula (see 3.5). ∎
@@ 5. The indecomposable condition
Let $\operatorname{\mathscr{C}}$ be a symmetric monoidal $\infty$-category with unit object $\mathbb{1}$. Recall that an idempotent of a commutative algebra $A\in\operatorname{CAlg}(\operatorname{\mathscr{C}})$ is by definition an idempotent element of the discrete ring $\pi_{0}(A):=\pi_{0}(\operatorname{Map}_{\operatorname{\mathscr{C}}}(\mathbb{1},A))$. We then say that $A$ is indecomposable if $A$ is nonzero and if it does not decompose as a product of two nonzero rings. This is equivalent to $\pi_{0}(A)$ being an indecomposable ring. See [NP24, Section 3] for more details.
@@
Let $\operatorname{\mathscr{C}}$ be a symmetric monoidal $\infty$-category and consider $R,S\in\operatorname{CAlg}(\operatorname{\mathscr{C}})$ with $S$ indecomposable. For any decomposition $R=\prod_{i=1}^{n}R_{i}$ into a finite product, we obtain
@@
Let $G$ be a finite group and consider $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. We say that $A$ satisfies the indecomposable condition (in short IC) if the commutative algebras $A^{hG}$ and $A^{tG}$ are indecomposable.
@@
Let $G$ be a finite group and let $A\in\operatorname{Fun}(BG,\operatorname{CAlg}^{\heartsuit})$ be a discrete commutative algebra with trivial $G$-action. Then $A$ satisfies IC if and only if $A$ and $A/|G|$ are indecomposable.
@@
$\pi_{0}(A^{hG})=H^{0}(G,A)=A\quad\mathrm{and}\quad\pi_{0}(A^{tG})=\widehat{H}^{0}(G,A)=A/|G|.$
@@
If $G$ is a nontrivial $p$-group, then the integers $\mathbb{Z}$ and any field $k$ of characteristic $p$ satisfy IC with respect to the trivial $G$-action.
@@
Let $G$ be a finite group and let $\mathbb{S}_{G}\in\operatorname{Fun}(BG,\operatorname{CAlg})$ be the sphere spectrum with trivial $G$-action. Then $\mathbb{S}_{G}$ satisfies IC if and only if $G$ is a nontrivial $p$-group for some prime number $p$.
@@
are indecomposable. Since the first quotient is $\mathbb{Z}$, and $[G]-|G|\in I$, we are reduced to checking when the two rings $\mathbb{Z}$ and $\mathbb{Z}/|G|$ are indecomposable. By the Chinese Remainder Theorem this holds if and only if $G$ is a nontrivial $p$-group for some prime $p$. ∎
@@ 6. The retraction condition
Let $G$ be a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. We say that $A$ satisfies the retraction condition (in short RC) if for all $H\subseteq G$ such that $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{tG}$, we have $H=G$.
@@
Suppose that $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfies RC. Then $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{hG}$ if and only if $H=G$. The “if” part is clear and for the “only if” part note that if $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{hG}$, then we can apply the canonical functor $\operatorname{Perf}(A)^{hG}\to\operatorname{Perf}(A)^{tG}$ to this retraction diagram to find, by RC, that $H=G$.
@@
Let $G$ be a finite group and let $X,Y$ be finite $G$-set. Suppose that the isotropy of $Y$ does not contain the trivial group. Let $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfy IC and RC. Then the canonical map
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Perf}(A)^{hG})}(A^{X},A^{Y})\xrightarrow{\sim}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Perf}(A)^{tG})}(A^{X},A^{Y})$
is an equivalence.
@@
We note that if $G=e$, then $Y=\varnothing$ and so $A^{Y}=0$. The canonical map in the theorem is then an equivalence as source and target are contractible. So without loss of generality we can assume that $G$ is nontrivial. Firstly, let us treat the special case in which $Y=G/G$ so that $A^{Y}=A$. Decomposing $X$ into its orbits $X\simeq\coprod_{i}G/H_{i}$, we see that $A^{X}\simeq\prod_{i}A^{G/H_{i}}$. Note that the unit $A$ is indecomposable both in $\operatorname{Perf}(A)^{hG}$ and $\operatorname{Perf}(A)^{tG}$ by the assumption IC for $A$. Remark 5.1 then implies that the relevant mapping space of commutative algebras in either category is given by
$\operatorname{Map}(A^{X},A)\simeq\coprod_{i}\operatorname{Map}(A^{G/H_{i}},A).$
We are therefore reduced to showing that passage to the quotient category induces an equivalence on the mapping spaces $\operatorname{Map}(A^{G/H},A)$. Indeed, we claim these mapping spaces are simultaneously empty or contractible, depending on $H$: If $H\not=G$, then we claim that both mapping spaces are empty as there are not even any unital maps from $A^{G/H}$ to $A$. This follows from the fact that $A$ satisfies RC and 6.2. If $H=G$, then both mapping spaces are contractible as the unit $A=A^{G/G}$ is the initial commutative algebra in both categories. We have therefore proved the theorem in the case that $Y=G/G$.
For the general case, decompose $Y$ into orbits $Y\simeq\coprod_{i}G/H_{i}$, and note that by our assumption $H_{i}\not=e$ for all $i$. Then in both categories we find that
@@
Let $G$ be a finite group and let $A\in\operatorname{CAlg}^{\heartsuit}$ be a $|G|$-torsion free discrete commutative ring in which no prime divisor of $|G|$ is invertible. Then $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ with trivial $G$-action satisfies RC.
@@
The naturality in $\mathscr{C}$ in 3.8 implies $B\otimes_{A}A^{G/H}\simeq B^{G/H}$ in $\operatorname{Perf}(B)^{tG}$. ∎

### v2

@@ Abstract.
Given a finite group $G$ and a commutative ring $G$-spectrum $R$, we study the separable commutative algebras in the category of compact $R$-modules. We isolate three conditions on the geometric fixed points of $R$ which ensure that every separable commutative algebra is standard, i.e. arises from a finite $G$-set. In particular we show that all separable commutative algebras in the categories of compact objects in $G$-spectra and in derived $G$-Mackey functors are standard provided that $G$ is a $p$-group. In these categories we also show that for a general finite group $G$, not all separable commutative algebras are standard. We finally discuss how the classification of separable commutative algebras in compact $G$-spectra varies if we require the existence of multiplicative norms. We show that if $G$ is solvable, then any separable commutative algebra therein that is normed is automatically standard. However, if $G$ is not solvable, we provide examples of separable commutative algebras that are normed but not standard.
@@ 1. Introduction
The study of separable commutative algebras in the setting of tensor-triangular geometry was initiated by Balmer in [BAL11]. Separable commutative algebras are useful in this context for several reasons. Firstly, they allow to form categories of modules internally to the world of triangulated categories, without the necessity of assuming that the triangulated category admits a model. Building on this, Balmer showed that restriction functors in (modular) representation theory can be understood as an extension-of-scalars functors [BAL15], and proved a generalization of the Neeman-Thomason Localization Theorem [BAL16b]. Secondly, separable commutative algebras allow us to formulate a going-up theorem in tt-geometry [BAL16a], which is a useful tool for calculating the Balmer spectrum of a triangulated category, and powerful descent results in tt-geometry [BAL16a],[BCH+24] and [BCH+25]. Finally, separable commutative algebras are tightly related to the étale topology of the given tt-category. A result of Neeman [NEE18] informally says that the only separable commutative algebras in the perfect derived category of a Noetherian scheme are those arising from finite étale morphisms. This connection with the étale topology was strengthened even further by the first two authors in [NP24] where they showed that separable commutative algebras recover finite étale algebras in various contexts where a notion of étaleness already exists.
@@
Informally this problem asks whether the étale topology which appears in modular representation theory is richer than what is produced by subgroups. A positive answer to this question was given by Balmer-Carlson [BC18] in the case of cyclic $p$-groups and by the first two authors [NP24] for $p$-groups of $p$-rank one. This question is completely open outside of these cases.
@@
    For all $1\not=U\subseteq W_{G}(K)$, the commutative algebras $(\Phi^{K}(R))^{hU}$ and $(\Phi^{K}(R))^{tU}$ are indecomposable.
@@
    If $\Phi^{K}(R)$ is a retract in $\operatorname{Mod}(\Phi^{K}(R))^{tU}$ of the coinduced algebra $(\Phi^{K}(R))^{U/V}$ for some $V\subseteq U\subseteq W_{G}(K)$, then $U=V$, see 6.1.
@@
  • (4)
    $\pi_{0}^{K}(R)$ is indecomposable and $\pi_{0}(\Phi^{K}R)$ is nonzero.
We will show that the above conditions are satisfied for all $K$ if $R=\mathbb{S}_{G}$ is the sphere $G$-spectrum, and also if $R=\operatorname{infl}_{e}^{G}\mathbb{Z}$, provided that $G$ is a $p$-group. The latter case has gained increasing interest for its connection to the category of derived $G$-Mackey functors of Kaledin, see [PSW22]. Our main result then states:
Theorem (9.8, 9.3).
@@
is an equivalence, and so any separable commutative algebra is standard. If moreover (4) holds for all $K\subseteq G$, then we have an equivalence
$R^{(-)}\colon\operatorname{Fin}^{\operatorname{op}}_{G}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega}).$
It follows that if $G$ is a $p$-group, then all separable commutative algebras are standard, both in compact $G$-spectra and in compact derived $G$-Mackey functors. In fact, for these categories, we can show that we have an equivalence before passing to groupoids. The techniques developed in this paper also allow us to produce some counterexamples. Indeed, we show in Section 10 that if $G=C_{6}$, then there are separable commutative algebras in the above two categories that are not standard.
As an immediate application of our work, we obtain the following computations of the Galois group in the sense of [MAT16].
Theorem (9.10, 10.10).
Let $p\neq q$ be prime numbers.
  • (a)
    If $G$ is a $p$-group, then the Galois groups of $\operatorname{Sp}_{G}$ is trivial.
  • (b)
    If $G$ is cyclic of order $pq$, then the Galois group of $\operatorname{Sp}_{G}$ is $\widehat{\mathbb{Z}}$.
In Section 11 we discuss how the classification of separable commutative algebras in compact $G$-spectra varies if we require the existence of multiplicative norms. We show that if $G$ is solvable, then any separable commutative algebra therein which is normed is automatically standard. However, if $G$ is not solvable, we provide examples of separable commutative algebras that are normed but not standard.
@@ Acknowledgements
We thank Bastiaan Cnossen, John Greenlees, Achim Krause and Phil Pützstück for useful conversations. We also thank the anonymous referee for their careful reading, helpful comments and for suggesting to consider the classification of normed separable algebras.
NN was supported by the SFB 1085 Higher Invariants in Regensburg. LP is grateful to Max Planck Institute for Mathematics in Bonn for its hospitality and financial support. LP was supported by the European Research Council (ERC) under Horizon Europe (grant No. 101042990) and by the SFB 1085 Higher Invariants in Regensburg. MR was supported by the Danish National Research Foundation through the Copenhagen Centre for Geometry and Topology (DNRF151) by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 427320536 - SFB 1442, as well as under Germany’s Excellence Strategy EXC 2044 390685587, Mathematics Münster: Dynamics-Geometry-Structure.
@@ 3. Homotopy $G$-fixed points and separable algebras
Given a group $G$ and $\mathscr{C}\in\operatorname{Fun}(BG,\operatorname{Cat}_{\infty})$, we can form
@@
where $K^{g}=gKg^{-1}$, and the restriction from $K$ to $H\cap K^{g}$ is along the conjugation map $H\cap K^{g}\subset K^{g}\cong K$. This is a consequence of [CLL25, Lemma 4.9] with $S=T=BG$ and $\mathcal{C}$ the $BG$-category given by
$(*)\,\,\,BG^{\mathrm{op}}\simeq BG\xrightarrow{A}\operatorname{CAlg}\xrightarrow{\operatorname{Mod}(-)}\operatorname{Cat}_{\infty}.$
@@
$\coprod_{g\in H\backslash G/K}G/(H\cap K^{g})$}BH{BH}BK{BK}BG{BG}
is an equivalence. One also checks that the limit-preserving extension of ()(*) to $\mathscr{S}_{/BG}$ sends $BH\to BG$ to $\operatorname{Mod}(A)^{hH}$.
@@
$\operatorname{Fin}_{G}^{\operatorname{op}}$$\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C}^{\operatorname{dual}})^{hG}$$\operatorname{Fin}_{H}^{\operatorname{op}}$$\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C}^{\operatorname{dual}})^{hH}$$\scriptstyle{(\mathbb{1}^{\bullet})^{hG}}$$\scriptstyle{\operatorname{res}^{G}_{H}}$$\scriptstyle{\operatorname{res}_{H}^{G}}$$\scriptstyle{(\mathbb{1}^{\bullet})^{hH}}$
@@ 4. Stable module categories and the Tate construction
$X{X}X{X}X/a{X/a}b\scriptstyle{b}$$\scriptstyle{ab}$$\scriptstyle{\mathrm{id}}$$\scriptstyle{a}$
@@
[…]
@@
[…]
@@
Since cofibres in $\operatorname{Cat}_{\infty}^{\mathrm{perf}}$ are given by the idempotent completion of the Verdier quotient, we need to check that the image of $\mathrm{Nm}_{G}$ is $\operatorname{thick}_{\otimes}\langle A^{G/e}\rangle$. By [CMN+24, Proposition 2.18] taking $\mathcal{A}=\operatorname{Perf}(A)\in\operatorname{Fun}(BG,\operatorname{Cat}_{\infty}^{\mathrm{perf}})$ there, we know that the image of the norm map agrees with $\operatorname{thick}(A^{G/e})$ which is automatically an ideal by an application of the projection formula (see 3.5). ∎
@@ 5. The indecomposable condition
Let $\operatorname{\mathscr{C}}$ be a stable presentably [^4] In fact, the reader can verify that additively symmetric monoidal and idempotent-complete is sufficient. symmetric monoidal $\infty$-category with unit object $\mathbb{1}$. Recall that an idempotent of a commutative algebra $A\in\operatorname{CAlg}(\operatorname{\mathscr{C}})$ is by definition an idempotent element of the discrete ring $\pi_{0}(A):=\pi_{0}(\operatorname{Map}_{\operatorname{\mathscr{C}}}(\mathbb{1},A))$. We then say that $A$ is indecomposable if $A$ is nonzero and if it does not decompose as a product of two nonzero rings. This is equivalent to $\pi_{0}(A)$ being an indecomposable ring. See [NP24, Section 3] for more details.
@@
Let $\operatorname{\mathscr{C}}$ be a stable presentably symmetric monoidal $\infty$-category and consider $R,S\in\operatorname{CAlg}(\operatorname{\mathscr{C}})$ with $S$ indecomposable. For any decomposition $R=\prod_{i=1}^{n}R_{i}$ into a finite product, we obtain
@@
Let $G$ be a finite group and consider $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. We say that $A$ satisfies the indecomposable condition (in short IC) if the commutative algebras $A^{hH}$ and $A^{tH}$ are indecomposable for all subgroups $1\not=H\subseteq G$.
@@
Let $G$ be a finite group and let $A\in\operatorname{Fun}(BG,\operatorname{CAlg}^{\heartsuit})$ be a discrete commutative algebra with trivial $G$-action. Then $A$ satisfies IC if and only if $A$ and $A/|H|$ are indecomposable for all $1\not=H\subseteq G$.
@@
$\pi_{0}(A^{hH})=H^{0}(H,A)=A\quad\mathrm{and}\quad\pi_{0}(A^{tH})=\widehat{H}^{0}(H,A)=A/|H|$
for all $H\subseteq G$. ∎
@@
If $G$ is a $p$-group, then the integers $\mathbb{Z}$ and any field $k$ of characteristic $p$ satisfy IC with respect to the trivial $G$-action.
@@
Let $G$ be a finite group and let $\mathbb{S}_{G}\in\operatorname{Fun}(BG,\operatorname{CAlg})$ be the sphere spectrum with trivial $G$-action. Then $\mathbb{S}_{G}$ satisfies IC if and only if $G$ is a $p$-group for some prime number $p$.
@@
are indecomposable. Since the first quotient is $\mathbb{Z}$, and $G]-|G|\in I$, we are reduced to checking when the two rings $\mathbb{Z}$ and $\mathbb{Z}/|G|$ are indecomposable. By the Chinese Remainder Theorem this holds if and only if $G$ is a nontrivial $p$-group for some prime $p$. It is only left to note that IC always holds if $G$ is the trivial group. ∎
@@ 6. The retraction condition
Let $G$ be a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. We say that $A$ satisfies the retraction condition for $G$ if for all $H\subseteq G$ such that $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{tG}$, we have $H=G$. We say that $A$ satisfies the retraction condition (in short RC) if for all subgroup $H\subseteq G$, the restricted commutative algebra $\operatorname{res}^{G}_{H}A$ satisfies the retraction condition for $H$.
@@
Suppose that $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfies the retraction condition for $G$. Then $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{hG}$ if and only if $H=G$. The “if” part is clear and for the “only if” part note that if $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{hG}$, then we can apply the canonical functor $\operatorname{Perf}(A)^{hG}\to\operatorname{Perf}(A)^{tG}$ to this retraction diagram to find, by RC, that $H=G$.
@@
Let $G$ be a finite group and let $X,Y$ be finite $G$-sets. Suppose that the isotropy of $Y$ does not contain the trivial group. Let $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfy IC and RC. Then both canonical maps
$\operatorname{Map}_{\operatorname{Fin}_{G}}(Y,X)\xrightarrow{\sim}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Perf}(A)^{hG})}(A^{X},A^{Y})\xrightarrow{\sim}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Perf}(A)^{tG})}(A^{X},A^{Y})$
are equivalences.
@@
We note that if $G=e$, then $Y=\varnothing$ and so $A^{Y}=0$. The canonical maps in the theorem are then equivalences as all mapping spaces involved are contractible. So without loss of generality we can assume that $G$ is nontrivial. Firstly, let us treat the special case in which $Y=G/G$ so that $A^{Y}=A$. Decomposing $X$ into its orbits $X\simeq\coprod_{i}G/H_{i}$, we see that $A^{X}\simeq\prod_{i}A^{G/H_{i}}$. Note that the unit $A$ is indecomposable both in $\operatorname{Perf}(A)^{hG}$ and $\operatorname{Perf}(A)^{tG}$ by the assumption IC for $A$. Remark 5.1 then implies that the relevant mapping space of commutative algebras in either category is given by
$\operatorname{Map}(A^{X},A)\simeq\coprod_{i}\operatorname{Map}(A^{G/H_{i}},A),$
and the analogous result holds obiously for $\operatorname{Fin}_{G}$. We are therefore reduced to showing that passage to the quotient category induces an equivalence on the mapping spaces $\operatorname{Map}(A^{G/H},A)$, and these are equivalent to $\operatorname{Map}_{\operatorname{Fin}_{G}}(*,G/H)$. Indeed, we claim these mapping spaces are simultaneously empty or contractible, depending on $H$: If $H\not=G$, then we claim that both mapping spaces are empty as there are not even any unital maps from $A^{G/H}$ to $A$. This follows from the fact that $A$ satisfies RC and 6.2. If $H=G$, then both mapping spaces are contractible as the unit $A=A^{G/G}$ is the initial commutative algebra in both categories. We have therefore proved the theorem in the case that $Y=G/G$.
For the general case, decompose $Y$ into orbits $Y\simeq\coprod_{i}G/H_{i}$, and note that by our assumption $H_{i}\not=e$ for all $i$. Then in all three categories we find that
@@
Let $G$ be a finite group and let $A\in\operatorname{CAlg}^{\heartsuit}$ be a $|G|$-torsion free discrete commutative ring in which no prime divisor of $|G|$ is invertible. Then $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ with trivial $G$-action satisfies the retraction condition for $G$. In fact $A$ satisfies RC.
@@
The naturality in $\mathscr{C}$ in 3.8 implies $B\otimes_{A}A^{H/K}\simeq B^{H/K}$ in $\operatorname{Perf}(B)^{tH}$ for any subgroups $K\subseteq H\subseteq G$. ∎
Example $6.$
Let $p$ be a prime number and $G$ be a finite $p$-group. It follows from Lemma 6.4 that $\mathbb{Z}$ satisfies RCRC. Using this and Lemma 6.5 we deduce that $\mathbb{S}$ with trivial $G$-action satisfies RCRC too.